OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + 2y = 4 x – 5y = 11

OpenStudy (anonymous):

solve first equation for x x=4-2y sub into second equation 4-2y-5y=11 (or it might look like this -2y+4-5y=11)